2.) Suppose the time X between calls to a phone center has a mean of 25 minutes and a standard deviation of 16 minutes. A simple random sample of 36 times between calls is taken. Find the probability that their average time is less than 20 minutes.
Answer #1

Solution :

Given that ,

mean =   = 25

standard deviation = = 16

n = 36

= 25

=  / n = 16/ 36=2.67

P( <20 ) = P[( - ) / < (20-25) /2.67 ]

= P(z <-1.87 )

Using z table  

= 0.0307   

probability=0.0307